# Non-Hermitian topology of transport in Chern insulators

### v1

@@ Abstract
Introduction — A physical system can have a non-Hermitian topological description if it is coupled to an environment and the resulting gains and losses are customized to reach a topological phase [1, 2]. Such a topological phase offers robust characteristics, with potential applications such as light funnels [3], amplifiers [4], and exponentially precise sensors [5, 6]. The experimental realization of these phases usually relies on platforms where gain and loss can be precisely controlled in each individual building block. Examples include ultracold atoms [7], optical systems [8, 9], and meta-materials [10, 11, 3, 12, 13, 14, 15].
A recent study takes an alternative approach, which leverages the quantum transport properties of a well-known Hermitian topological phase – the quantum Hall phase in a GaAs/AlGaAs heterostructure [16]. In the presence of multiple current sources, the current-voltage characteristics of the quantum Hall phase show robust features that originate from the non-Hermitian topology of the conductance matrix, not of the Hamiltonian. The chiral edge modes ensure that charge propagates between neighboring contacts only in one direction, meaning that the multi-terminal conductance matrix becomes numerically equivalent to the Hamiltonian matrix of the Hatano-Nelson (HN) chain [17], a prototypical model of non-Hermitian topology. In effect, each terminal plays the role of a ‘site’ in the HN chain, while the edge modes propagating between adjacent contacts play the role of nearest-neighbor, one-way ‘hoppings.’
@@
To illustrate this concept, we turn to graphene, one of the most accessible and highly-tunable platforms for realizing quantum Hall physics. Using a toy model, we first highlight the correspondence between chiral edge modes and the non-Hermitian topology of the conductance matrix. This correspondence remains valid even as the chemical potential is swept through the charge neutrality point and the edge modes change direction. Next, we introduce inhomogeneities into the multi-terminal conductance matrix, showing that they cause the emergence of an additional phase.
Transport in graphene and its connection to non-Hermitian topology — We use a toy model for graphene consisting of spinless fermions hopping on a honeycomb lattice. The Hamiltonian takes the form
@@
[…]
We introduce a perpendicular magnetic field by changing the vertical hoppings in Fig. 1(a) from $t$ to $t\exp(i\phi n_{x}^{{\phantom{{\dagger}}}})$, where $\phi/2\pi$ is the flux per hexagon, in units of the flux quantum. This gaps out the Dirac cones and results in the formation of Landau levels, symmetrically distributed around the charge neutrality point, $E=0$ [18] (see the supplemental material (SM) for details [19]). Since the model is spinless, sweeping the chemical potential from $\mu>0$ to $\mu<0$ reverses the edge mode propagation direction from counter-clockwise to clockwise, leading to a direct transition between phases with Chern numbers $+1$ and $-1$.
Figure 1: Panel (a): Graphene lattice with two sites (shown as white and black circles) in each unit cell (marked by the yellow dashed contour). $\vec{e}_{x}^{{\phantom{{\dagger}}}}$ and $\vec{e}_{y}^{{\phantom{{\dagger}}}}$ are the Bravais vectors and $t$ is the nearest neighbor hopping. Panel (b): Schematic of a graphene disk with six contacts (blue) attached to its boundaries and labeled from one to six. The transmission between adjacent leads in the clockwise and anticlockwise directions is represented by red and green arrows, respectively. Panel (c): Conductance matrix $G$ for a disk of graphene with six leads [as in (b)], deep in a quantum Hall phase with Chern number $-1$ ($\mu/t=-0.2$, $\phi=0.18$). The conductance (in units of $e^{2}/h$) is quantized between adjacent leads in the downstream direction of the edge modes and is zero in the opposite direction. The full conductance matrix is equivalent to a Hatano-Nelson Hamiltonian under PBC whereas the sub-matrix (enclosed by blue dashed lines) is equivalent to the same under OBC. Panel (d): SPD plotted for the $G$ matrix in panel (c) against the site index (lead index) on a log-linear scale under OBC (non-Hermitian skin effect) and PBC.
Using the Kwant package [20], we attach $N$ equally-spaced ideal leads to the boundary of a finite-sized, disk-shaped system, as shown in Fig. 1(b) for $N=6$. Numerical details are described in the SM [19]. This models an experimental setup in which $N$ contacts are connected to a graphene sample. In the linear regime, the currents injected into each of the contacts, $\vec{I}=(I_{1},\ldots,I_{N})$, are related to the potentials on each of the contacts, $\vec{V}=(V_{1},\ldots,V_{N})$, as $\vec{I}=G\vec{V}$, where $G$ is the $N\times N$ conductance matrix. At zero-temperature and in the zero bias voltage limit, the elements of $G$ take the form
@@
For the six-terminal system shown in Fig. 1(b), setting $\mu=-0.2t$ and $\phi=0.18$, such that the system is deep in the quantum Hall phase with Chern number $-1$, leads to the conductance matrix shown in Fig. 1(c). This matrix is identical to the non-Hermitian Hamiltonian matrix of the Hatano-Nelson tight-binding model [17], a chain of orbitals connected by nonreciprocal nearest-neighbor hoppings. Thus, each of the diagonal elements $G_{ii}$ can be thought of as being analogous to the onsite energy of a site in the fictitious chain. The off-diagonal elements are analogous to hopping integrals connecting the different fictitious sites. In this case, each of the sites has an onsite energy equal to $+1$, one-way hoppings equal to $-1$ connect the $j^{\rm th}$ and $(j-1)^{\rm th}$ sites, and since $G_{61}=-1$, the chain has periodic boundary conditions (PBC). Realizing an analogue of the Hatano-Nelson model with open boundary conditions (OBC) can be done by grounding the $6^{\rm th}$ contact (setting the potential $V_{6}=0$). In this case the currents and voltages corresponding to the first five terminals would be related by the $5\times 5$ conductance sub-matrix, $G_{\rm OBC}$, shown by the dashed box in Fig. 1(c).
@@
Inhomogeneity and topological phase transitions — We are interested in studying the topology of the conductance matrix away from the quantum Hall plateau regime, when bulk transport is significant and all elements of $G$ are nonzero. To this end, we follow Refs. [23, 16], and focus only on the OBC setup in the following. The $N^{\rm th}$ contact is grounded at all times and only the $(N-1)\times(N-1)$ sub-block of the conductance matrix, $G_{\rm OBC}$, is considered. We use a topological invariant, $w_{\rm PD}^{{\phantom{{\dagger}}}}$, that can be determined directly from this conductance matrix, and which predicts the existence of a non-Hermitian skin effect. To compute it, we first subtract the average of the diagonal elements from the conductance matrix, $\hat{G}=G_{\rm OBC}-\lambda\mathds{1}$, where $\lambda={\rm tr}(G_{\rm OBC})/(N-1)$ and $\mathds{1}$ is the identity matrix. Writing the polar decomposition of this matrix, $\hat{G}=QP$, where $Q$ is unitary and $P$ is a positive-definite matrix, yields
@@
where $X={\rm diag}(1,2,\ldots,N-1)$ encodes the lead index, and $\mathcal{T}$ is the trace per unit volume evaluated over the middle indices, far from the ends (see SM for details [19]).
As expected, we confirm that $w_{\rm PD}^{{\phantom{{\dagger}}}}$ correctly predicts the appearance of a non-Hermitian skin effect in the conductance matrix of Fig. 1(d). Deep in the quantum Hall plateau regime, where the boundary-localized chiral modes produce one-way nearest-neighbor hoppings of the fictitious Hatano-Nelson chain, the invariant in Eq. (3) is locked to the Chern number. We have verified that sweeping the chemical potential across the charge neutrality point leads to a direct transition in the Chern number from $+1$ to $-1$, and simultaneously to a direct transition of $w_{\rm PD}^{{\phantom{{\dagger}}}}$, from $-1$ to $+1$. Correspondingly, there is a change of direction in the non-Hermitian skin effect: all eigenvectors of the conductance matrix are all localized either on the left or the right side of the chain, depending on the sign of $\mu$.
In any experiment, we expect that the conductance matrix will be inhomogeneous due to multiple factors. These include the varying quality of different contacts, unequal distances between them, as well as disorder and defects in the sample. For simplicity, we obtain inhomogeneous conductance matrices $G$ by randomly changing the potential on each site of the honeycomb lattice, $\mu\rightarrow\mu+\delta_{i}^{{\phantom{{\dagger}}}}$, where $\delta_{i}^{{\phantom{{\dagger}}}}$ is chosen from a uniform distribution $(-\delta,\delta)$.
@@
for $n$ independent realizations of the random onsite potential, each producing a different conductance matrix $G_{i}$. Figure 2 shows that for $\delta=0$ the conductance matrix exhibits a direct transition from $w_{\rm PD}^{{\phantom{{\dagger}}}}=+1$ to $-1$ as the chemical potential is swept across the charge neutrality point, as mentioned above. Turning on inhomogeneities, however, leads to the appearance of a trivial phase of the conductance matrix, $\langle w_{\rm PD}^{{\phantom{{\dagger}}}}\rangle=0$, for which the current-voltage characteristics do not show the exponential profile of the non-Hermitian skin effect (see SM for details [19]). We observe that increasing $\delta$ leads to an enlargement of this phase, which appears for a wider range of chemical potential $\mu$.
Figure 2: $\langle w_{\rm PD}^{{\phantom{{\dagger}}}}\rangle$ evaluated for conductance matrices of a monolayer spinless graphene disk with 10 leads, one of which is grounded (9 lead OBC matrix) as a function of varying $\delta/t$ and chemical potential $\mu/t$. The trivial phase ($\langle w_{\rm PD}^{{\phantom{{\dagger}}}}\rangle=0$) becomes more prominent close to the charge neutrality point $\mu/t=0$ as $\delta$ increases. Here, $\phi=0.18$.
@@
The intervening trivial non-Hermitian phase of the conductance matrix appears in a region where the bulk of the graphene is conducting, due to the fact that random potential fluctuations broaden the zeroth Landau level. However, it is not caused by the presence of conducting bulk modes. To show this, we consider a different averaging scheme, where we first average the conductance matrix over the random realizations of onsite disorder, and only afterwards compute the topological invariant:
$\langle w_{\rm PD}^{{\phantom{{\dagger}}}}\rangle_{2}^{{\phantom{{\dagger}}}}=w_{\rm PD}^{{\phantom{{\dagger}}}}\left(\frac{1}{n}\sum_{i=1}^{n}\hat{G_{i}}\right).$ (5)
@@
Figure 3: Panel (a): Absolute value of disorder-averaged conductances $\langle G\rangle_{1,2}$ (purple) and $\langle G\rangle_{2,1}$ (green), showing the transition between two quantum Hall phases with Chern numbers $C=1$ and $C=-1$, for a spinless graphene disk with $N=40$ contacts. In these phases, the conductance through the bulk is negligible, as indicated by the $\langle G\rangle_{1,20}\times 100$ (multiplied by a factor of 100 for ease of visibility) (blue). This bulk conductance becomes non-zero in the transition region between the two quantum Hall phases, indicating the diffusive transport regime [see SM for details [19]]. Panel (b): evaluated $\langle w_{\rm PD}^{{\phantom{{\dagger}}}}\rangle$ shows the presence of a trivial phase close to $\mu/t=0$ between two non-Hermitian topological phases ($+1$ and $-1$). This is in contrast to $\langle w_{\rm PD}^{{\phantom{{\dagger}}}}\rangle_{2}$ which shows a direct transition between the two phases. Here, $\delta/t=1.2$, and $\phi=0.18$. Panel (c): A $6\times 6$ conductance sub-matrix at $\mu=-0.04t$ for a single realization of onsite disorder, indicating the inhomogeneities in its entries. Panel (d): A $6\times 6$ sub-matrix of the disorder-averaged conductance matrix $\langle G\rangle$ at $\mu=-0.04t$ exhibiting the uniformity in its entries (see SM for details [19]).
The true reason for emergence of the extra phase between $\langle w_{\rm PD}^{{\phantom{{\dagger}}}}\rangle=\pm 1$ phases is due to inhomogeneities in the conductance matrix and is a natural consequence of the current conservation rules. Since the sum of all elements of $G$ over any column must vanish due to Kirchhoff’s laws, $\sum_{i=1}^{L}G_{ij}^{{\phantom{{\dagger}}}}=0$, random conductance fluctuations between different pairs of leads necessarily lead to random changes on the conductance matrix diagonal. On the level of the fictitious, 1D Hatano-Nelson chain simulated by $G$, this corresponds to the presence of onsite disorder, which is known to lead to an extra $\langle w_{\rm PD}^{{\phantom{{\dagger}}}}\rangle=0$ phase (see Ref. [23] and SM [19]).
Conclusion — We have theoretically investigated the non-Hermitian topological behavior associated with multiple-current-source transport in Chern insulators, using the quantum Hall phases of a honeycomb graphene model as an example. Our simulations demonstrated the presence and robustness of non-Hermitian phases in the conductance matrix of this system, which can easily be tuned through chemical potential variations. This variation can be achieved in practice via gate voltage adjustments, making graphene a promising platform to realize current-voltage relations protected non-Hermitian topology.

### v2

# Non-Hermitian topology of transport in the quantum Hall phases in graphene

@@ Abstract
I Introduction
A physical system can have a non-Hermitian topological description if it is coupled to an environment and the resulting gains and losses are customized to reach a topological phase [1, 2]. Such a topological phase offers robust characteristics, with potential applications such as light funnels [3], amplifiers [4], and exponentially precise sensors [5, 6]. The experimental realization of these phases usually relies on platforms where gain and loss can be precisely controlled in each individual building block. Examples include ultracold atoms [7], optical systems [8, 9], and meta-materials [10, 11, 3, 12, 13, 14, 15].
A recent study takes an alternative approach, which leverages the quantum transport properties of a well-known Hermitian topological phase – the quantum Hall phase in a GaAs/AlGaAs heterostructure [16]. It is based on the insight that in the limits of maximum non-reciprocity, the Hamiltonian for a Hatano-Nelson (HN) chain [17] accurately describes the long-time dynamics of the chiral edge states in a quantum Hall system [18].
In the presence of multiple current sources, the current-voltage characteristics of the quantum Hall phase show robust features that originate from the non-Hermitian topology of the conductance matrix, not of the Hamiltonian. The chiral edge modes ensure that charge propagates between neighboring contacts only in one direction, meaning that the multi-terminal conductance matrix becomes numerically equivalent to the Hamiltonian matrix of the HN chain, a prototypical model of non-Hermitian topology. In effect, each terminal plays the role of a ‘site’ in the HN chain, while the edge modes propagating between adjacent contacts play the role of nearest-neighbor, one-way ‘hoppings.’
@@ I Introduction
To illustrate this concept, we turn to graphene, one of the most accessible and highly-tunable platforms for realizing quantum Hall physics. In the GaAs/AlGaAs heterostructure based 2DEG used in the previous study [16], the direction of edge state propagation could be reversed by reversing the direction of the magnetic field, which would involve going through a classical Hall regime. A graphene based experimental setup would offer more tunability, as the direction of the edge states can additionally be reversed through the application of a gate voltage.
Using a toy model, we first highlight the correspondence between chiral edge modes and the non-Hermitian topology of the conductance matrix. This correspondence remains valid even as the chemical potential is swept through the charge neutrality point and the edge modes change direction. Next, we introduce inhomogeneities into the multi-terminal conductance matrix, showing that they cause the emergence of an additional phase. This additional non-Hermitian trivial phase corresponds to the vanishing of the exponential profile of current and voltage vectors, and this occurs in a regime where graphene has finite non-reciprocal conductance between its terminals, i.e., the non-Hermitian trivial phase of the conductance matrix cannot be traced back to a trivial insulating phase in graphene.
II Transport in graphene and its connection to non-Hermitian topology —
We use a toy model for graphene consisting of spinless fermions hopping on a honeycomb lattice. The Hamiltonian takes the form
@@ II Transport in graphene and its connection to non-Hermitian topology —
where $\mu$ represents the on-site energy of an electron at the site index $i$, $c^{{\phantom{{\dagger}}}}_{i}$ $(c^{\dagger}_{i})$ represent its annihilation (creation) operator, $\langle\dots\rangle$ denotes nearest neighbors, and $t$ is the hopping parameter. The unit cells are labeled by integers $(n_{x}^{{\phantom{{\dagger}}}},n_{y}^{{\phantom{{\dagger}}}})$ in the $\vec{e}_{x}^{{\phantom{{\dagger}}}},\vec{e}_{y}^{{\phantom{{\dagger}}}}$ directions, respectively [see Fig. 1(a)].
We introduce a perpendicular magnetic field by changing the vertical hoppings in Fig. 1(a) from $t$ to $t\exp(i\phi n_{x}^{{\phantom{{\dagger}}}})$, where $\phi/2\pi$ is the flux per hexagon, in units of the flux quantum. This gaps out the Dirac cones and results in the formation of Landau levels, symmetrically distributed around the charge neutrality point, $E=0$ [19] (see Appendix A for details). Since the model is spinless, sweeping the chemical potential from $\mu>0$ to $\mu<0$ reverses the edge mode propagation direction from counter-clockwise to clockwise, leading to a direct transition between phases with Chern numbers +1+1 and 1-1.
Figure 1: Panel (a): Graphene lattice with two sites (shown as white and black circles) in each unit cell (marked by the yellow dashed contour). $\vec{e}_{x}^{{\phantom{{\dagger}}}}$ and $\vec{e}_{y}^{{\phantom{{\dagger}}}}$ are the Bravais vectors and $t$ is the nearest neighbor hopping. Panel (b): Schematic of a graphene disk with six contacts (blue) attached to its boundaries and labeled from one to six. The transmission between adjacent leads in the clockwise and anticlockwise directions is represented by red and green arrows, respectively. Panel (c): Conductance matrix $G$ for a disk of graphene (radius $r=40$ in units of the lattice constant) with six leads [as in (b)], deep in a quantum Hall phase with Chern number 1-1 ($\mu/t=-0.2$, $\phi=0.18$). The conductance (in units of $e^{2}/h$) is quantized between adjacent leads in the downstream direction of the edge modes and is zero in the opposite direction. The full conductance matrix is equivalent to a Hatano-Nelson Hamiltonian under PBC whereas the sub-matrix (enclosed by blue dashed lines) is equivalent to the same under OBC. Panel (d): SPD plotted for the $G$ matrix in panel (c) against the site index (lead index) on a log-linear scale under OBC (non-Hermitian skin effect) and PBC.
Using the Kwant package [20], we attach $N$ equally-spaced ideal leads to the boundary of a finite-sized, disk-shaped system, as shown in Fig. 1(b) for $N=6$. Numerical details are described in Appendix B. This models an experimental setup in which $N$ contacts are connected to a graphene sample. In the linear regime, the currents injected into each of the contacts, $\vec{I}=(I_{1},\ldots,I_{N})$, are related to the potentials on each of the contacts, $\vec{V}=(V_{1},\ldots,V_{N})$, as $\vec{I}=G\vec{V}$, where $G$ is the $N\times N$ conductance matrix. At zero-temperature and in the zero bias voltage limit, the elements of $G$ take the form
@@
For the six-terminal system shown in Fig. 1(b), setting $\mu=-0.2t$ and $\phi=0.18$, such that the system is deep in the quantum Hall phase with Chern number 1-1, leads to the conductance matrix shown in Fig. 1(c). This matrix is identical to the non-Hermitian Hamiltonian matrix of the Hatano-Nelson tight-binding model [17], a chain of orbitals connected by nonreciprocal nearest-neighbor hoppings. Thus, each of the diagonal elements $G_{ii}$ can be thought of as being analogous to the onsite energy of a site in the fictitious chain. The off-diagonal elements are analogous to hopping integrals connecting the different fictitious sites. In this case, each of the sites has an onsite energy equal to +1+1, one-way hoppings equal to 1-1 connect the $j^{\rm th}$ and $(j-1)^{\rm th}$ sites, and since $G_{61}=-1$, the chain has periodic boundary conditions (PBC). Realizing an analogue of the Hatano-Nelson model with open boundary conditions (OBC) can be done by grounding the $6^{\rm th}$ contact (setting the potential $V_{6}=0$). In this case the currents and voltages corresponding to the first five terminals would be related by the $5\times 5$ conductance sub-matrix, $G_{\rm OBC}$, shown by the dashed box in Fig. 1(c).
@@
III Inhomogeneity and topological phase transitions —
We are interested in studying the topology of the conductance matrix away from the quantum Hall plateau regime, when bulk transport is significant and all elements of $G$ are nonzero. To this end, we follow Refs. [23, 16], and focus only on the OBC setup in the following. The $N^{\rm th}$ contact is grounded at all times and only the $(N-1)\times(N-1)$ sub-block of the conductance matrix, $G_{\rm OBC}$, is considered. We use a topological invariant, $w_{\rm PD}^{{\phantom{{\dagger}}}}$, that can be determined directly from this conductance matrix, and which predicts the existence of a non-Hermitian skin effect. To compute it, we first subtract the average of the diagonal elements from the conductance matrix, $\hat{G}=G_{\rm OBC}-\lambda\mathds{1}$, where $\lambda={\rm tr}(G_{\rm OBC})/(N-1)$ and $\mathds{1}$ is the identity matrix. Writing the polar decomposition of this matrix, $\hat{G}=QP$, where $Q$ is unitary and $P$ is a positive-definite matrix, yields
@@ III Inhomogeneity and topological phase transitions —
where $X={\rm diag}(1,2,\ldots,N-1)$ encodes the lead index, and $\mathcal{T}$ is the trace per unit volume evaluated over the middle indices, far from the ends (see Appendix B for details).
Figure 2: $\langle w_{\rm PD}^{{\phantom{{\dagger}}}}\rangle$ evaluated for conductance matrices of a monolayer spinless graphene disk ($r=60$) with 10 leads, one of which is grounded (9 lead OBC matrix) as a function of varying $\delta/t$ and chemical potential $\mu/t$. The trivial phase ($\langle w_{\rm PD}^{{\phantom{{\dagger}}}}\rangle=0$) becomes more prominent close to the charge neutrality point $\mu/t=0$ as $\delta$ increases. Here, $\phi=0.18$.
As expected, we confirm that $w_{\rm PD}^{{\phantom{{\dagger}}}}$ correctly predicts the appearance of a non-Hermitian skin effect in the conductance matrix of Fig. 1(d). Deep in the quantum Hall plateau regime, where the boundary-localized chiral modes produce one-way nearest-neighbor hoppings of the fictitious Hatano-Nelson chain, the invariant in Eq. (3) is locked to the Chern number. We have verified that sweeping the chemical potential across the charge neutrality point leads to a direct transition in the Chern number from +1+1 to 1-1, and simultaneously to a direct transition of $w_{\rm PD}^{{\phantom{{\dagger}}}}$, from 1-1 to +1+1. Correspondingly, there is a change of direction in the non-Hermitian skin effect: all eigenvectors of the conductance matrix are all localized either on the left or the right side of the chain, depending on the sign of $\mu$.
In any experiment, we expect that the conductance matrix will be inhomogeneous due to multiple factors. These include the varying quality of different contacts, unequal distances between them, disorder and defects in the sample, as well as sample-specific quantum fluctuation effects (universal conductance fluctuations) [24].
For simplicity, we obtain inhomogeneous conductance matrices $G$ by randomly changing the potential on each site of the honeycomb lattice, $\mu\rightarrow\mu+\delta_{i}^{{\phantom{{\dagger}}}}$, where $\delta_{i}^{{\phantom{{\dagger}}}}$ is chosen from a uniform distribution $(-\delta,\delta)$.
@@
for $n$ independent realizations of the random onsite potential, each producing a different conductance matrix $G_{i}$.
Figure 3: Panel (a): Disorder-averaged Chern number ($C$) for a monolayer spinless graphene disk (r=15) plotted as a function of the disorder strength $\delta/t$ and chemical potential $\mu/t$. Panel (b): Horizontal line cuts from panel (a) showing the disorder-averaged Chern number exhibiting a direct transition, i.e. varying continuously between $\pm 1$ for various non-zero disorder strengths $\delta/t$, without any intermediate trivial phase. Here, $\phi=0.18,\kappa=0.01$.
Figure 4: Panel (a): Absolute value of disorder-averaged conductances $\langle G\rangle_{1,2}$ (purple) and $\langle G\rangle_{2,1}$ (green), showing the transition between two quantum Hall phases with Chern numbers $C=1$ and C=1C=-1, for a spinless graphene disk ($r=240$) with $N=40$ contacts. In these phases, the conductance through the bulk is negligible, as indicated by the $\langle G\rangle_{1,20}\times 100$ (multiplied by a factor of 100 for ease of visibility) (blue). This bulk conductance becomes non-zero in the transition region between the two quantum Hall phases, indicating the diffusive transport regime [see Appendix D]. Panel (b): evaluated $\langle w_{\rm PD}^{{\phantom{{\dagger}}}}\rangle$ shows the presence of a trivial phase close to $\mu/t=0$ between two non-Hermitian topological phases (+1+1 and 1-1). This is in contrast to $\langle w_{\rm PD}^{{\phantom{{\dagger}}}}\rangle_{2}$ which shows a direct transition between the two phases. Here, $\delta/t=1.2$, and $\phi=0.18$. Panel (c): A $6\times 6$ conductance sub-matrix at $\mu=-0.04t$ for a single realization of onsite disorder, indicating the inhomogeneities in its entries. Panel (d): A $6\times 6$ sub-matrix of the disorder-averaged conductance matrix $\langle G\rangle$ at $\mu=-0.04t$ exhibiting the uniformity in its entries (see Appendix C).
@@
Figure 2 shows that for $\delta=0$ the conductance matrix exhibits a direct transition from $w_{\rm PD}^{{\phantom{{\dagger}}}}=+1$ to 1-1 as the chemical potential is swept across the charge neutrality point, as mentioned above. Turning on inhomogeneities, however, leads to the appearance of a trivial phase of the conductance matrix, $\langle w_{\rm PD}^{{\phantom{{\dagger}}}}\rangle=0$, for which the current-voltage characteristics do not show the exponential profile of the non-Hermitian skin effect (see Appendix C). We observe that increasing $\delta$ leads to an enlargement of this phase, which appears for a wider range of chemical potential $\mu$.
The intervening trivial non-Hermitian phase of the conductance matrix appears in a region where the graphene disk is not a trivial insulator. We confirm this by computing a local Chern marker for the graphene disk, known as the Localizer index. This is a topological invariant evaluated using real-space methods and is equivalent to the Chern number $C$ for quantum Hall systems [25, 26]. In particular, using the position operators $\bar{X}$ and $\bar{Y}$ of the graphene disk encoding the position information of the lattice sites and the Hamiltonian $H$, a composite Hermitian operator known as the spectral Localizer $L$ is constructed as:
$L(\bar{X},\bar{Y},H)=\kappa[\sigma_{x}\otimes\bar{X}+\sigma_{y}\otimes\bar{Y}]+\sigma_{z}\otimes H,$ (5)
where $\kappa$ is a scaling factor chosen appropriately to ensure compatible weights for $H,\bar{X},$ and $\bar{Y}$ [25, 26]. The topological index is then computed as:
$C=\frac{1}{2}\text{\rm sig}[L(\bar{X},\bar{Y},H)],$ (6)
where sig refers to the signature of a matrix (number of positive eigenvalues minus the number of negative eigenvalues). Fig. 3 shows that for non-zero disorder strength $\delta$, the topological index (equivalent to the Chern number $C$) exhibits the absence of a trivial insulating phase with zero Chern number, i.e. it varies continuously between the values $\pm 1$ without having a plateau at zero.
In the region where the non-Hermitian trivial phase develops, the bulk of the graphene is conducting, due to the fact that random potential fluctuations broaden the zeroth Landau level. This is shown in Fig. 4 (a) as a non-zero conductance value across the graphene disk $G_{1,20}$ for a disk with 40 contacts and also in Appendix D. In the thermodynamic limit, the transport through the bulk will vanish for all chemical potentials except at the quantum Hall plateau transition $\mu=0$. On the level of the effective Hatano-Nelson chain realized by the multi-terminal conductance matrix, the conductance between adjacent leads mimics the “hopping” between adjacent “sites” of the chain. Thus, in order to meaningfully go towards the thermodynamic limit of the effective Hatano-Nelson chain we have to scale up both the radius of the graphene disk as well as the total number of contacts, in such a way that the distance between neighboring contacts remains constant. We implement this finite-size scaling study by increasing the radius $r$ of the graphene disk while simultaneously increasing the number of contacts $N$ attached to its boundaries ($N=r/6$), grounding 1 lead every time to get the OBC matrix. The corresponding $\langle w_{\rm PD}^{{\phantom{{\dagger}}}}\rangle$ averaged over 1000 disordered configurations shows that the extra non-Hermitian phase gets better resolved as the size of the original graphene disk increases (see Fig. 5(a)). In another scaling study, we fix the size of the graphene disk ($r=300$) and change the number of contacts used in the calculation of the conductance matrix. We start by attaching 50 leads to the graphene disk and ground $p$ out of these to give a $(50-p)\times(50-p)$ conductance matrix under OBC. The scaling behavior of the $\langle w_{\rm PD}^{{\phantom{{\dagger}}}}\rangle$ (averaged over 1500 independent onsite disorder realizations) is shown in Fig. 5(b) for different choices of $p$ corresponding to different sizes $N$ of the OBC conductance matrix, again indicating that the trivial phase is better resolved as $N$ increases. Finally, we emphasize that the equivalence between the topology of a d-dimensional system and its (d-1) dimensional edge [18] could be obtained in our setting when the conductance matrix has contributions solely from the edge. This would correspond to the case where the radius of the graphene disk is made infinitely large without increasing the number of contacts. The separation between the contacts in this case will be much larger than the localization length of the bulk states, resulting in a complete suppression of bulk transmission.
Figure 5: Finite-size scaling: (a) $\langle w_{\rm PD}^{{\phantom{{\dagger}}}}\rangle$ for a disk with increasing radius $r$ and proportionately increasing number of leads $N=r/6$ showing that the trivial phase gets better resolved as the size of the graphene disk increases. (b) $\langle w_{\rm PD}^{{\phantom{{\dagger}}}}\rangle$ for different sizes $N$ of the OBC conductance matrix $G_{\rm OBC}^{{\phantom{{\dagger}}}}$ as a function of the chemical potential $\mu$ indicating the presence of the trivial phase for different number of leads. Here, $\delta=1.2t$.
@@
The presence of conducting bulk modes is not the cause of the the non-Hermitian trivial phase. To show this, we consider a different averaging scheme, where we first average the conductance matrix over the random realizations of onsite disorder, and only afterwards compute the topological invariant:
[…]
@@
The true reason for emergence of the extra phase between $\langle w_{\rm PD}^{{\phantom{{\dagger}}}}\rangle=\pm 1$ phases is due to inhomogeneities in the conductance matrix and is a natural consequence of the current conservation rules. Since the sum of all elements of $G$ over any column must vanish due to Kirchhoff’s laws, $\sum_{i=1}^{L}G_{ij}^{{\phantom{{\dagger}}}}=0$, random conductance fluctuations between different pairs of leads necessarily lead to random changes on the conductance matrix diagonal. On the level of the fictitious, 1D Hatano-Nelson chain simulated by $G$, this corresponds to the presence of onsite disorder, which is known to lead to an extra $\langle w_{\rm PD}^{{\phantom{{\dagger}}}}\rangle=0$ phase (see Ref. [23] and Appendix E).
IV Conclusion
We have theoretically investigated the non-Hermitian topological behavior associated with multiple-current-source transport in Chern insulators, using the quantum Hall phases of a honeycomb graphene model as an example. Our simulations demonstrated the presence and robustness of non-Hermitian phases in the conductance matrix of this system, which can easily be tuned through chemical potential variations. This variation can be achieved in practice via gate voltage adjustments, making graphene a promising platform to realize current-voltage relations protected non-Hermitian topology.